OpenStudy (anonymous):

can someone help me with 4 questions ??? :(

OpenStudy (anonymous):

I assure you that you are much more likely to be held once you actually post your questions. @dianamathematics

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

are you guys there

OpenStudy (anonymous):

Note that I =Prt. In each choice under "Term", you are given 'I' which is the simple interest, the time period in years which is 't', and the interest rate 'r'. For each just multiply those values given to see with which one of A,B,C,D the answer and then put the corresponding value in. I'll do the first one for you: "$72 over 3 years at an interest rate of 3%. Here my t = 3, r = 0.03 (you have to divide the percent by 100 to get the decimal number) and my 'I' which is the simple interest equals $72. Now plug everything in to I = Prt and solve for 'P':\[\bf 72=P(0.03)(3) \implies P = \frac{ 72 }{ 0.03(3) }=$800\]This is the same as option B.) hence the first one is B.). Now solve the remaining 2 in a similar manner :). @dianamathematics